OpenStudy (anonymous):

you have two complementary angles. if the measure of the larger angle is 15 degrees more than 2 times the measure of the smaller angle, find the measures of the 2 angles.

OpenStudy (anonymous):

The key is to express both angles in terms of only a single variable. Also key is knowing the definition of complementary angles.

OpenStudy (anonymous):

i dont know how to create the two equations..can u show me...

OpenStudy (anonymous):

i got x + y = 180 and 3x-28+x= 180

OpenStudy (anonymous):

Why = 180?

OpenStudy (anonymous):

oh nevr mind..diferent problm sorry

OpenStudy (anonymous):

i got x + y=90 and 2x +15 +y=90

OpenStudy (anonymous):

is that correct?

OpenStudy (anonymous):

You're right that x + y = 90. Now, replace one of those variables so you only have one variable to deal with. You can solve from there.

OpenStudy (anonymous):

so u mean it would be 2x + 15 +x=90

OpenStudy (anonymous):

???

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

thnkz